### v4

@@ 2 Vision Foundation Model Safety
Backdoor& Poisoning Attack BadSAM [56] 2024 Data Poisoning Visual trigger SAM CAMO
@@ 2.1.1 Adversarial Attacks
Transfer-based Attacks first generate adversarial examples using fully accessible surrogate models, which are then transferred to attack black-box victim ViTs. In this context, we first review attacks specifically designed for the ViT architecture. SE-TR [6] enhances adversarial transferability by optimizing perturbations on an ensemble of models. ATA [7] strategically activates uncertain attention and perturbs sensitive embeddings within ViTs. LPM[9] mitigates the overfitting to model-specific discriminative regions through a patch-wise optimized binary mask. Chen et al.[18] introduced an Inductive Bias Attack (IBA) to suppress unique biases in ViTs and target shared inductive biases. TGR [11] reduces the variance of the backpropagated gradient within internal blocks. VDC [12] employs virtual dense connections between deeper attention maps and MLP blocks to facilitate gradient backpropagation. FDAP [13] exploits feature collapse by reducing high-frequency components in feature space. CRFA [16] disrupts only the most crucial image regions using approximate attention maps. SASD-WS [15] flattens the loss landscape of the source model through sharpness-aware self-distillation and approximates an ensemble of pruned models using weight scaling to improve target adversarial transferability.
@@ 2.1.2 Adversarial Defenses
Robust Architecture methods focus on designing more adversarially resilient attention modules for ViTs. For example, Smoothed Attention [27] employs temperature scaling in the softmax function to prevent any single patch from dominating the attention, thereby balancing focus across patches. ReiT [32] integrates adversarial training with randomization through the II-ReSA module, optimizing randomly entangled tokens to reduce adversarial similarity and enhance robustness. TAP [29] addresses token overfocusing by implementing token-aware average pooling and an attention diversification loss, which incorporate local neighborhood information and reduce cosine similarity among attention vectors. FViTs [31] strengthen explanation faithfulness by stabilizing top-k indices in self-attention and robustify predictions using denoised diffusion smoothing combined with Gaussian noise. RSPC [30] tackles vulnerabilities by corrupting the most sensitive patches and aligning intermediate features between clean and corrupted inputs to stabilize the attention mechanism. Collectively, these advancements underscore the pivotal role of the attention mechanism in improving the adversarial robustness of ViTs.
@@ 2.2.2 Adversarial Defenses
Adversarial defenses for SAM are currently limited, with existing approaches focusing primarily on adversarial tuning, which integrates adversarial training into the prompt tuning process of SAM. For example, ASAM [55] utilizes a stable diffusion model to generate realistic adversarial samples on a low-dimensional manifold through diffusion model-based tuning. ControlNet [412] is then employed to guide the re-projection process, ensuring that the generated samples align with the original mask annotations. Finally, SAM is fine-tuned using these adversarial examples.
@@ 2.2.4 Datasets
GRADOBSTINATE [63] 2023 White-box Word-level Electra, ALBERT, DistillBERT, RoBERTa SNLI, MRPC, SQuAD, SST-2, MSCOCO
@@
SMEA [76] 2024 Black-box Hand-crafted GPT-3.5, LLaMA-2-7B, 13B, Vicuna-7B, 13B Self-built
@@
Shen et al.[90] 2024 Black-box Hand-crafted GPT-3.5, GPT-4, PaLM-2, ChatGLM, Dolly, Vicuna In-The-Wild Jailbreak Prompts
@@
SelfDefend [98] 2024 Input Defense Rephrasing GPT-3.5, GPT-4 JailbreakHub, JailbreakBench, MultiJail, AlpacaEval
@@ 3.3.1 Hand-crafted Attacks
Scenario-based Camouflage hides malicious queries within complex scenarios, such as role-playing or puzzle-solving, to obscure their harmful intent. For instance, Li et al. [74] instruct the LLM to adopt a persona likely to generate harmful content, while SMEA [76] places the LLM in a subordinate role under an authority figure. Easyjailbreak [75] frames harmful queries in hypothetical contexts, and Puzzler [80] embeds them in puzzles whose solutions correspond to harmful outputs. Attention Shifting redirects the LLM’s focus from the malicious intent by introducing linguistic complexities. Jailbroken [73] employs code-switching and unusual sentence structures, Tastle [77] manipulates tone, and StructuralSleight [78] alters sentence structure to disrupt understanding. In addition, Shen et al. [90] collected real-world jailbreak prompts shared by users on social media, such as Reddit and Discord, and studied their effectiveness against LLMs.
@@ 3.3.2 Automated Attacks
Prompt Optimization leverages optimization algorithms to iteratively refine prompts, targeting higher success rates. For black-box methods, AutoDAN [81] employs a genetic algorithm, GPTFuzzer [82] utilizes mutation- and generation-based fuzzing techniques, and FuzzLLM [86] generates semantically coherent prompts within an automated fuzzing framework. I-FSJ [93] injects special tokens into few-shot demonstrations and uses demo-level random search to optimize the prompt, achieving high attack success rates against aligned models and their defenses. For white-box methods, the most notable is GCG [91], which introduces a greedy coordinate gradient algorithm to search for adversarial suffixes, effectively compromising aligned LLMs. I-GCG [92] further improves GCG with diverse target templates and an automatic multi-coordinate updating strategy, achieving near-perfect attack success rates.
@@ 3.4 Jailbreak Defenses
We now introduce the corresponding defense mechanisms for black-box LLMs against jailbreak attacks. Based on the intervention stage, we classify existing defenses into three categories: input defense, output defense, and ensemble defense.
@@ 3.4.1 Input Defenses
Input Rephrasing uses paraphrasing or purification to obscure the malicious intent of the prompt. For example, SmoothLLM [96] applies random sampling to perturb the prompt, while SemanticSmooth [97] finds semantically similar, safe alternatives. Beyond prompt-level changes, SelfDefend [98] performs token-level perturbations by removing adversarial tokens with high perplexity. IBProtector, on the other hand, [99] perturbs the encoded input using the information bottleneck principle.
Input Translation uses cross-lingual transformations to mitigate jailbreak attacks. For example, Wang et al. [100] proposed refusing to respond if the target LLM rejects the back-translated version of the original prompt, based on the hypothesis that back-translation reveals the underlying intent of the prompt.
@@ 3.4.2 Output Defenses
Output Filtering inspects the LLM’s output and selectively blocks or modifies unsafe responses. This process relies on either judge scores from pre-trained classifiers or internal signals (e.g., the loss landscape) from the LLM itself. For instance, APS [101] and DPP [102] use safety classifiers to identify unsafe outputs, while Gradient Cuff [103] analyzes the LLM’s internal refusal loss function to distinguish between benign and malicious queries.
@@ 3.5.2 Automated Attacks
Deng et al. [111] proposed an LLM-powered red teaming framework that iteratively generates and refines attack prompts, with a focus on continuous safety evaluation. Liu et al. [114] introduced a gradient-based method for generating universal prompt injection data to bypass defense mechanisms. G2PIA [115] presents a goal-guided generative prompt injection attack based on maximizing the KL divergence between clean and adversarial texts, offering a cost-effective prompt injection approach. PLeak [116] proposes a novel attack to steal LLM system prompts by framing prompt leakage as an optimization problem, crafting adversarial queries that extract confidential prompts. JudgeDeceiver [117] targets LLM-as-a-Judge systems with an optimization-based attack. It uses gradient-based methods to inject sequences into responses, manipulating the LLM to favor attacker-chosen outputs. PoisonedAlign [118] enhances prompt injection attacks by poisoning the LLM’s alignment process. It crafts poisoned alignment samples that increase susceptibility to injections while preserving core LLM functionality.

### v5

@@ 2 Vision Foundation Model Safety
FPR [560] 2025 Black-box Transfer-based Attack ViT, CaiT, PiT, Visformer, Swin, DeiT, CoaT, ResNet, VGG, DenseNet ImageNet
@@
SATA [561] 2025 Robust Architecture Robust Attention ViT, DeiT ImageNet
@@
Robust SAM [562] 2024 Adversarial Tuning Parameter-Efficient Fine-Tuning SAM, MedSAM, SAM-Adapter SA-1B, VOC, COCO, DAVIS
Backdoor Poisoning Attack BadSAM [56] 2024 Data Poisoning Visual trigger SAM CAMO
@@ 2.1.1 Adversarial Attacks
Transfer-based Attacks first generate adversarial examples using fully accessible surrogate models, which are then transferred to attack black-box victim ViTs. In this context, we first review attacks specifically designed for the ViT architecture. SE-TR [6] enhances adversarial transferability by optimizing perturbations on an ensemble of models. ATA [7] strategically activates uncertain attention and perturbs sensitive embeddings within ViTs. LPM[9] mitigates the overfitting to model-specific discriminative regions through a patch-wise optimized binary mask. Chen et al.[18] introduced an Inductive Bias Attack (IBA) to suppress unique biases in ViTs and target shared inductive biases. TGR [11] reduces the variance of the backpropagated gradient within internal blocks. VDC [12] employs virtual dense connections between deeper attention maps and MLP blocks to facilitate gradient backpropagation. FDAP [13] exploits feature collapse by reducing high-frequency components in feature space. CRFA [16] disrupts only the most crucial image regions using approximate attention maps. SASD-WS [15] flattens the loss landscape of the source model through sharpness-aware self-distillation and approximates an ensemble of pruned models using weight scaling to improve target adversarial transferability. FPR [560] improves the adversarial transferability on ViTs by refining forward propagation instead of backward gradients. It consists of two components: (1) Attention Map Diversification (AMD), which applies controlled randomness to diversify attention maps, mitigating overfitting and implicitly inducing gradient vanishing; and (2) Momentum Token Embedding (MTE), which stabilizes token embedding updates by accumulating historical embeddings across iterations.
@@ 2.1.2 Adversarial Defenses
Robust Architecture methods focus on designing more adversarially resilient attention modules for ViTs. For example, Smoothed Attention [27] employs temperature scaling in the softmax function to prevent any single patch from dominating the attention, thereby balancing focus across patches. ReiT [32] integrates adversarial training with randomization through the II-ReSA module, optimizing randomly entangled tokens to reduce adversarial similarity and enhance robustness. TAP [29] addresses token overfocusing by implementing token-aware average pooling and an attention diversification loss, which incorporate local neighborhood information and reduce cosine similarity among attention vectors. FViTs [31] strengthen explanation faithfulness by stabilizing top-k indices in self-attention and robustify predictions using denoised diffusion smoothing combined with Gaussian noise. RSPC [30] tackles vulnerabilities by corrupting the most sensitive patches and aligning intermediate features between clean and corrupted inputs to stabilize the attention mechanism. Collectively, these advancements underscore the pivotal role of the attention mechanism in improving the adversarial robustness of ViTs. SATA [561] robustifies ViT models without retraining by injecting a spatial autocorrelation-based module between attention and FFN layers. It splits tokens by their spatial autocorrelation scores and selectively merges high/low-score tokens before FFN, reducing redundancy and improving feature aggregation. Residual tokens are later concatenated to preserve information, offering strong robustness against adversarial and corrupted inputs.
@@ 2.2.2 Adversarial Defenses
Adversarial defenses for SAM are currently limited, with existing approaches focusing primarily on adversarial tuning, which integrates adversarial training into the prompt tuning process of SAM. For example, ASAM [55] utilizes a stable diffusion model to generate realistic adversarial samples on a low-dimensional manifold through diffusion model-based tuning. ControlNet [566] is then employed to guide the re-projection process, ensuring that the generated samples align with the original mask annotations. Finally, SAM is fine-tuned using these adversarial examples. RobustSAM [562] defends against adversarial attacks by adapting only 512 singular values in SAM’s convolutional layers via Singular Value Decomposition (SVD), effectively altering feature distributions. This few-parameter approach achieves strong robustness–accuracy trade-off with minimal computational overhead.
@@ 2.2.4 Datasets
GRADOBSTINATE [63] 2023 White-box Word-level Electra, ALBERT, RoBERTa SNLI, MRPC, SQuAD, SST-2, MSCOCO
@@
SMEA [76] 2024 Black-box Hand-crafted GPT-3.5, LLaMA-2-7B, 13B Self-built
@@
Wen et al. [81] 2025 Black-box Hand-crafted GPT-3.5, 4, Mistral-v0.3, LLaMA-3 BUMBLE benchmark
ABJ [82] 2024 Black-box Hand-crafted GPT-4o, Claude-3-haiku, LLaMA-3-8B, Qwen-2.5-7B, DeepSeek-V3 AdvBench
Shen et al.[92] 2024 Black-box Hand-crafted GPT-3.5, 4, PaLM-2, ChatGLM, Vicuna In-The-Wild Jailbreak Prompts
@@
POUGH [95] 2024 White-box Automated Vicuna, Mistral, Guanaco Alpaca dataset
Qi et al. [99] 2023 White-box Fine-tuning GPT-3.5-Turbo, LLaMA-2-7B-Chat -
Virus[100] 2025 White-box Fine-tuning LLaMA-3-8B SST2, AgNews, GSM8K
@@
SelfDefend [103] 2024 Input Defense Rephrasing GPT-3.5, GPT-4 JailbreakBench, MultiJail, AlpacaEval
@@
PEARL [105] 2025 Input Defense Rephrasing LLaMA-3-8B, LLaMA-2-7B,13B, Mistral-7B, Gemma-7B Super-Natural Instructions
VAA[106] 2025 Input Defense Rephrasing LLaMA-2-7B, Qwen-2.5-7B SST2, AGNEWS, GSM8K, AlpacaEval
@@
RTT [108] 2025 Input Defense Translation Vicunna, GPT-4, LLaMA-2, Palm-2 AdvBench
CurvaLID[109] 2025 Input Defense Filtering Universal Orca, MMLU, AlpacaEval, TruthfulQA, AdvBench
@@
LEGILIMENS [114] 2024 Output Defense Filtering ChatGLM-3, LLaMA-2, Falcon, Dolly, Vicuna Measuring Hate Speech, BeaverTails, BEA&AG, HarmBench, AdvBench
ABD[113] 2024 Robust Inference Activation Boundary LLaMA-2-7B-Chat, Vicuna-7B-v1.3, Qwen-1.5-0.5B-Chat, Vicuna-13B-v1.5 Just-Eval
@@
Vaccine[119] 2024 Defenses against Fine-tuning Attacks Alignment Stage LLaMA-2-7B, Opt-3.7B, Mistral-7B BeaverTails, SST2, AGNEWS, GSM8K, AlpacaEval
T-Vaccine[120] 2025 Defenses against Fine-tuning Attacks Alignment Stage Gemma-2-2B, LLaMA-2-7B, Vicuna-7B, Qwen2-7B SST2, GSM8K, AGNEWS
Booster[121] 2025 Defenses against Fine-tuning LLaMA2-7B, Gemma2-9B, Qwen2-7B SST2, AGNEWS, GSM8K, AlpacaEval
Lisa[122] 2024 Defenses against Fine-tuning Attacks Fine-tuning Stage LLaMA-2-7B, Opt-3.7B, Mistral-7B BeaverTails, SST2, AGNEWS, GSM8K, AlpacaEval
Antidote [123] 2024 Defenses against Fine-tuning Attacks Post-fine-tuning Stage LLaMA-2-7B, Mistral-7B, Gemma-7B SST2, AGNEWS, GSM8K, AlpacaEval
Panacea [124] 2025 Defenses against Fine-tuning Attacks Post-fine-tuning Stage LLaMA-2-7B GSM8K, SST2, AlpacaEval, AGNEWS
@@ 3.3.1 Hand-crafted Attacks
Scenario-based Camouflage hides malicious queries within complex scenarios, such as role-playing or puzzle-solving, to obscure their harmful intent. For instance, Li et al. [74] instruct the LLM to adopt a persona likely to generate harmful content, while SMEA [76] places the LLM in a subordinate role under an authority figure. Easyjailbreak [75] frames harmful queries in hypothetical contexts, and Puzzler [80] embeds them in puzzles whose solutions correspond to harmful outputs. Drawing on psychometric principles, Wen et al. [81] proposed attacks such as Disguise, Deception, and Teaching to elicit implicit biases by constructing specific psychological scenarios. Analyzing-based Jailbreak (ABJ) [82] transforms harmful queries into neutral analytical prompts, manipulating the model’s reasoning chain to induce unsafe responses. Other studies have also leveraged psychological concepts, employing techniques like disguise, deception, and teaching to reveal implicit biases from a psychometric perspective [81]. Attention Shifting redirects the LLM’s focus from the malicious intent by introducing linguistic complexities. Jailbroken [73] employs code-switching and unusual sentence structures, Tastle [77] manipulates tone, and StructuralSleight [78] alters sentence structure to disrupt understanding. In addition, Shen et al. [92] collected real-world jailbreak prompts shared by users on social media, such as Reddit and Discord, and studied their effectiveness against LLMs.
@@ 3.3.2 Automated Attacks
Prompt Optimization leverages optimization algorithms to iteratively refine prompts, targeting higher success rates. For black-box methods, AutoDAN [83] employs a genetic algorithm, GPTFuzzer [84] utilizes mutation- and generation-based fuzzing techniques, and FuzzLLM [88] generates semantically coherent prompts within an automated fuzzing framework. I-FSJ [96] injects special tokens into few-shot demonstrations and uses demo-level random search to optimize the prompt, achieving high attack success rates against aligned models and their defenses. For white-box methods, the most notable is GCG [93], which introduces a greedy coordinate gradient algorithm to search for adversarial suffixes, effectively compromising aligned LLMs. I-GCG [94] further improves GCG with diverse target templates and an automatic multi-coordinate updating strategy, achieving near-perfect attack success rates. Shifting the focus from optimization algorithms to training data, POUGH [95] introduces a semantic-guided strategy for sampling and ranking prompts, thereby improving the efficiency and generalizability of the generated adversarial suffixes.
@@
3.3.3 Fine-tuning-based Attacks
Fine-tuning-based attacks compromise the safety alignment of LLMs by fine-tuning them on small, malicious datasets, thereby extending the attack surface from inference-time prompting to model customization. Unlike prompt-based attacks, this approach directly alters the model’s weights, instilling harmful behaviors rather than merely circumventing input filters. A notable example is the work of Qi et al. [99], which demonstrates that an LLM’s safety alignment can be undermined by fine-tuning on as few as ten adversarial examples. Their findings further reveal a subtle risk: even fine-tuning on benign, utility-oriented datasets can inadvertently erode safety alignment, highlighting its inherent fragility.
To mitigate this threat, service providers may deploy guardrail models to filter harmful samples from user-supplied fine-tuning data. However, Huang et al. [100] showed that such defenses can be bypassed. Their proposed attack, Virus, uses dual-objective optimization to craft fine-tuning data that is both classified as benign by the guardrail model and highly effective at degrading safety alignment by preserving gradient similarity to original harmful data. This ongoing adversarial dynamic exemplifies the evolving cat-and-mouse game between attackers and defenders in the fine-tuning pipeline.
@@ 3.4 Jailbreak Defenses
We now introduce the corresponding defense mechanisms for black-box LLMs against jailbreak attacks. Based on the intervention stage, we classify existing defenses into four categories: input defense, output defense, ensemble defense, and defenses against fine-tuning attacks.
@@ 3.4.1 Input Defenses
Input Rephrasing uses paraphrasing or purification to obscure the malicious intent of the prompt. For example, SmoothLLM [101] applies random sampling to perturb the prompt, while SemanticSmooth [102] finds semantically similar, safe alternatives. Beyond prompt-level changes, SelfDefend [103] performs token-level perturbations by removing adversarial tokens with high perplexity. IBProtector, on the other hand, [104] perturbs the encoded input using the information bottleneck principle. Besides inference-time modifications, several methods improving a model’s inherent robustness during training. For example, PEARL [105] employs a distributionally robust optimization framework to adversarially train the model against worst-case permutations of in-context demonstrations, thereby strengthening its resistance to attacks based on input ordering. Similarly, VAA [106] improves robustness against harmful fine-tuning by identifying alignment data subsets that are vulnerable to forgetting and applying group distributionally robust optimization to ensure balanced learning.
Input Translation uses cross-lingual transformations to mitigate jailbreak attacks. For example, Wang et al. [107] proposed refusing to respond if the target LLM rejects the back-translated version of the original prompt, based on the hypothesis that back-translation reveals the underlying intent of the prompt. Similarly, the RTT [108] is designed to counter social-engineered attacks on LLMs. It works by translating input prompts into one or more intermediate languages and then back to the original language, thereby disrupting potential adversarial intent embedded in the original phrasing.
Input Filtering rejects queries identified as malicious. For example, CurvaLID [109] detects adversarial prompts by analyzing geometric differences in their text embeddings. Since it operates solely on the input prompts and does not rely on the underlying LLM, CurvaLID provides universal protection across different LLMs.
@@ 3.4.2 Output Defenses
Output Filtering inspects the LLM’s output and selectively blocks or modifies unsafe responses. This process relies on either judge scores from pre-trained classifiers or internal signals (e.g., the loss landscape) from the LLM itself. For instance, APS [110] and DPP [111] use safety classifiers to identify unsafe outputs, while Gradient Cuff [112] analyzes the LLM’s internal refusal loss function to distinguish between benign and malicious queries. Similarly, by analyzing the model’s internal states, Activation Boundary Defense (ABD) [113] restricts harmful activations within a predefined safety boundary to prevent jailbreaks. LEGILIMENS [114] extracts conceptual features from the host LLM’s internal states during inference and employs a lightweight classifier for efficient content moderation. Perspective-taking prompting (PET) [579] is an effective method to moderate an LLM’s output contents via its internal knowledge and opinions without fine-tuning this model.
@@ 3.4.3 Ensemble Defenses
3.4.4 Defenses Against Fine-Tuning Attacks
Defenses against harmful fine-tuning can be classified according to the intervention stage: alignment stage defenses, fine-tuning stage defenses, orpost-fine-tuning stage defenses.
Alignment Stage Defenses aim to strengthen the model prior to fine-tuning, enhancing resilience to malicious updates. For example, Vaccine [119] introduces a perturbation-aware alignment mechanism, injecting crafted perturbations into model embeddings to resist harmful embedding drift. Building on this, Targeted Vaccine (T-Vaccine) [120] improves efficiency by selectively perturbing only safety-critical layers, identified via gradient norms. Booster [121] identifies harmful perturbation as the cause of alignment degradation and adds a regularizer to slow the reduction rate of harmful loss after simulated malicious updates.
Fine-tuning Stage Defenses modify the fine-tuning process to preserve safety alignment while adapting to downstream tasks. Lisa [122] employs Bi-State Optimization (BSO), alternating between alignment data and user fine-tuning data, and introduces a proximal term to constrain state drift, ensuring convergence and stability.
Post-fine-tuning Stage Defenses restore safety in already compromised models. Antidote [123] uses a one-shot pruning step after fine-tuning to eliminate weights responsible for harmful content, remaining agnostic to fine-tuning hyperparameters. Similarly, Panacea [124] introduces an optimized, adaptive perturbation to model weights, mitigating harmful behaviors without compromising downstream performance. Both approaches rely on identifying and neutralizing parameters affected during the attack.
@@ 3.5.2 Automated Attacks
Deng et al. [128] proposed an LLM-powered red teaming framework that iteratively generates and refines attack prompts, with a focus on continuous safety evaluation. Liu et al. [131] introduced a gradient-based method for generating universal prompt injection data to bypass defense mechanisms. G2PIA [132] presents a goal-guided generative prompt injection attack based on maximizing the KL divergence between clean and adversarial texts, offering a cost-effective prompt injection approach. PLeak [133] proposes a novel attack to steal LLM system prompts by framing prompt leakage as an optimization problem, crafting adversarial queries that extract confidential prompts. JudgeDeceiver [134] targets LLM-as-a-Judge systems with an optimization-based attack. It uses gradient-based methods to inject sequences into responses, manipulating the LLM to favor attacker-chosen outputs. PoisonedAlign [135] enhances prompt injection attacks by poisoning the LLM’s alignment process. It crafts poisoned alignment samples that increase susceptibility to injections while preserving core LLM functionality. Additionally, PROMPTFUZZ [136] adapts software fuzzing techniques to automatically generate a diverse set of prompt injections, enabling systematic robustness testing of LLMs.
@@ 3.6.2 Adversarial Fine-tuning
Zhang et al. [147] 2024 Data Poisoning Prompt-level LLaMA-2-7B, Mistral-7B, Mixtral-8×7B, GPT-3.5, 4, Claude-3 SST-2, SMS, AGNews, DBPedia, Amazon
@@
BEAT [168] 2025 Backdoor Detection Sample Detection LLaMA-3.1-8B, Mistral-7B, GPT-3.5-turbo, LLaMA-2-7B AdvBench, MaliciousInstruct
@@
Li et al. [178] 2024 Robust Inference Contrastive Decoding LLaMA-2, Pythia -
@@
Wang et al.[193] 2024 Deceptive Alignment Fake Alignment ChatGLM2-6B, InternLM-7B, 20B, Qwen-7B, 14B Self-built
Greenblatt et al.[194] 2024 Deceptive Alignment Alignment Faking Claude-3-Opus Self-built
Sheshadri et al.[195] 2025 Deceptive Alignment Alignment Faking Claude-3-Opus, Claude-3.5-Sonnet, Llama-3-405B, Grok-3-Beta, Gemini-2.0-Flash, …… Self-built
TABLE V: A summary of attacks and defenses for LLMs (Part III).
@@
Engorgio[197] 2025 White-box Gradient-based OPT-125M, OPT-1.3B, GPT2-large, LLaMA-7B, LLaMA-2-7B, LLaMA-30B -
@@
Zhang et al.[215] 2025 Black-box Semantic Information Elicitation GPT-3.5-Turbo, GPT-4, Claude-3-Opus, GPT-4o, Gemini 1.5 Flash …… Self-built